OpenStudy (anonymous):

In a certain county, the number of Charter Schools is eight less than twice the number of Alternative Schools. We know that there are 36 Charter Schools in the county. How many Alternative Schools are in the county?

OpenStudy (ajprincess):

Let the number of Charter Schools be x and number of Alternative schools be y. x=36 x=2y-8 36=2y-8 Can u find y @bancoocoo1?

OpenStudy (ajprincess):

Let the number of Charter Schools be x and number of Alternative schools be y. Twice the number of Alternative Schools=2y. Eight less than twice the number of Alternative Schools=2y-8 less than means u need to subtract. Charter Schools is equal to eight less than twice the number of Alternative Schools That is x=2y-8 Getting this @bancoocoo1?

OpenStudy (anonymous):

MOSTLY

OpenStudy (anonymous):

I THINK I KIND OF AM STATING TO GET IT

OpenStudy (ajprincess):

if it is still confusing plz ask befor I sign out:)

OpenStudy (anonymous):

OK YA IT IS I THINK THAT I HAVE ONE MORE EXAMPLE I NEED U TO HELP WITH Currently, Sam's coach is three times older than Sam. When Sam started soccer 5 years ago, his coach was five times his age. How old is Sam now?

OpenStudy (anonymous):

I WISH I UNDERSTOOD THIS AS EASILY AS U DO

OpenStudy (ajprincess):

Can u jst try to make the eqn? then I can find where u r stuck. It will be easy for me to explain u.:)

OpenStudy (anonymous):

FIRST I DONT KNOW WHO TO MAKE AS A VARIABLE SAM OR THE CHOCH OR BOTH

OpenStudy (ajprincess):

First start making both as variables.

OpenStudy (anonymous):

BUT MAYEBY IT IS 3S=5S-5 S IS SAM

OpenStudy (anonymous):

OK SO IT IS (5S-5)+3S I REALLY DONT GET IT AT ALL

OpenStudy (anonymous):

COULD U EXPLAIN YOUR WHOLOE THOUGHT PROCESS OR SOMETHING I NEED TO KNOW HOW TO DO THIS

OpenStudy (ajprincess):

Let us take sentence by sentence and try to get the info from them. Let us assume Sam's current age to be s and coach's current age be c. Currently, Sam's coach is three times older than Sam means c=3s Following this?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

UR REALLY GOOD AT THIS WHAT GRADE ARE YOU IN IM IN 6TH

OpenStudy (ajprincess):

Nw u need to find the age of them before 5 yrs ago. Can u tell me the age of coach before 5 yrs in terms of c? Well I am in uni nw. I've finished schooling.

OpenStudy (anonymous):

O AND I THINK IT MIGHT BE S5=C

OpenStudy (ajprincess):

hmm no.

OpenStudy (ajprincess):

let us say ur age is 11. What was ur age before 5 yrs?

OpenStudy (anonymous):

OK SO IT IS C-5=S5

OpenStudy (anonymous):

BE FOR 5 YEARS WHAST DOES THAT MEAN MINUS 5?

OpenStudy (ajprincess):

yup.

OpenStudy (anonymous):

OK WELL I WOULD BE 6

OpenStudy (anonymous):

SO IT WOULD BE S3+S2=C THEN

OpenStudy (ajprincess):

So u subtracted 5 from 11 to find 6. Similarly to find the coach's and Sam's ages before 5 yrs u hav to subtract 5 from their current ages.

OpenStudy (ajprincess):

Find the coach's age and Sam's age before 5 yrs seperately and tell me.

OpenStudy (anonymous):

C-5=S3-5

OpenStudy (anonymous):

RIGHT?

OpenStudy (ajprincess):

What u have written is right. coach's age before 5 yrs=c-5 sam's age before 5 yrs=s-5 Before 5 yrs coach's age was 5 times the age of Sam. So c-5=5(s-5) Getting it?

OpenStudy (anonymous):

THAT IS THE BEFORE 5 YEARS EQUATION OR THE WHOLE EQUATION I DON'T UNDERSTAND THAT EQUATION IM SORRY NOT THAT ADVANCED IN MATH

OpenStudy (ajprincess):

Before 5 yrs c-5=5(s-5) Currently c=3s Getting it nw?

OpenStudy (anonymous):

OK YA SO BEFORE 5 YEARS IT IS SAM'S CURRENT AGE MINUS 5 TIMES 5 EQUALS COACHES CURRENT AGE MINUS 5? RIGHT?

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

BUT THAT IS RIGHT SO THE EQUATION FOR NOW IS WHAT?

OpenStudy (anonymous):

MY TEST IS IN AN HOUR AND I NEED TO UNDERSTAND

OpenStudy (ajprincess):

Nw u hav two equations with two variables. Using them u hav to find s.

OpenStudy (anonymous):

HOW DO U MAKE THEM ONE EQUATION?

OpenStudy (ajprincess):

c-5=5(s-5) c=3s In ur first eqn u can put 3s in place of c. Then u will get 3s-5=5(s-5) Getting it?

OpenStudy (anonymous):

SO THAT IS THE WHOLE EQUATION I THINK I GET IT BUT I KNOW U HAVE TO GO SO I WILL FIGURE IT OUT

OpenStudy (ajprincess):

yup. if u solve for s then u r done:)

OpenStudy (anonymous):

THANKS SO MUCH I HOPE I DO WELL ON THE TEST WISH ME GOOD LUCK